1. Two runners decide to race. Runner A is faster than Runner B, so Runner A starts 13.5 meters behind the starting line and Runner B starts on the starting line. Both runners start at the same time. Runner A runs at a speed of 9.65 m/s and Runner B runs at a speed of 8.40 m/s. How far down the track from the starting line does Runner A catch Runner B.
1
Expert's answer
2020-12-14T12:15:06-0500

"x_b=v_{0b}t"


"x_a=x_0+x_b=v_{0\u0430}t"


So, we have


"x_b=\frac{x_0}{v_{0a}/v_{0b}-1}=\frac{13.5}{9.65/8.40-1}=90.72(m)"